OpenStudy (anonymous):

What is the slope of a line that is perpendicular to the line whose equation is ax+by=c? A) a/b B) b/a C) -b/a D) c/b

OpenStudy (anonymous):

condition of two line is perpendicular is multiplication of their slopes is=-1.

OpenStudy (nincompoop):

perpendicular to the slope so two slopes must be the negative reciprocal of each other an example would be 3/4 and -4/3 or -2 = 1/2

OpenStudy (anonymous):

So, A?

OpenStudy (anonymous):

are you know the slope of given line?

OpenStudy (anonymous):

I don't know whether it's A or B

OpenStudy (nincompoop):

ax+by=c in slope-intercept form it will be y = (-ax+c)/b we the have -a/b so its negative reciprocal - (b/-a) = b/a

OpenStudy (anonymous):

Thx

OpenStudy (nincompoop):

@hartnn @ganeshie8 please check my response if logically valid.

hartnn (hartnn):

yes, thats correct :)

OpenStudy (nincompoop):

thank you